OpenStudy (anonymous):

help in verifying identities:

OpenStudy (anonymous):

\[\sin ^{2}x \tan ^{2}x +\sec ^{2}x=\sec ^{4}x\]

zepdrix (zepdrix):

Hello Miss KatKat! c: Do you remember how to convert tangent into sines and cosines?

OpenStudy (anonymous):

Yep!

zepdrix (zepdrix):

\[\Large\rm \sin ^{2}x \left(\color{orangered}{\tan^{2}x}\right) +\sec ^{2}x=\sec ^{4}x\]

zepdrix (zepdrix):

So what can we do with that orange guy? :d

OpenStudy (anonymous):

\[\sin ^{2}x \div \cos ^{2}x\]

zepdrix (zepdrix):

\[\Large\rm \sin ^{2}x \left(\color{orangered}{\frac{\sin^2x}{\cos^2x}}\right) +\sec ^{2}x=\sec ^{4}x\]Mmm k good good good. From there what can we do? Oooo hold up, I misread the problem >.< Lemme think about this one a sec.. not sure if this is the way to go.

OpenStudy (anonymous):

haha how are you making those equations look like that? also, this problem probably uses the power reducing formulas.

zepdrix (zepdrix):

If you right-click my fancy math stuff, `Show Math As` > `TeX Commands` You can see the way I typed it in. Gotta know a few fun little tricks to be able to do it hehe :D

zepdrix (zepdrix):

Power reductions for secant? Hmm I know the power reduction for secant using integrals. And I also know the Half-Angle formula helps to reduce powers, but that messes up our angle x. What formula you speaking of? :U

zepdrix (zepdrix):

brb i need some food in my belly :U

zepdrix (zepdrix):

Are you sure you copied the problem down correctly? This one seems kinda crazy 0_o

OpenStudy (anonymous):

shoot i didnt. \[\Large\rm \sec ^{2}x \left(\color{magenta}{\tan^{2}x}\right) +\sec ^{2}x=\sec ^{4}x\]

OpenStudy (anonymous):

thats it, i apologize

zepdrix (zepdrix):

Still a 4th power though? :U

OpenStudy (anonymous):

theses are those formulas i was talking about

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

could this work?

OpenStudy (anonymous):

one sec

zepdrix (zepdrix):

Yah those are the Half-Angle Identities, but notice how the `u` changes to `2u`? That would mess up our angle :) So we can't use those. darn.

zepdrix (zepdrix):

Remember your Pythagorean Identity involving tangent and secant? That one should help us :)

OpenStudy (anonymous):

i feel like i could still use those? could we split the 4th power into sec^{2}x squared?

zepdrix (zepdrix):

Yes. But we don't need to. What we want to do is recall our Pythagorean Identity for tangent/secant.\[\Large\rm \tan^2x+1=\sec^2x\]That one look familiar? We want to solve it for tan^2x.

OpenStudy (anonymous):

\[\Large\rm \sec ^{2}x \left(\color{orangered}{\sec^{2}x-1}\right) +\sec ^{2}x=\sec ^{4}x\]

OpenStudy (anonymous):

oh i got it!

zepdrix (zepdrix):

Ooo nice job \c:/ hehe getting the hang of those colors too huh?

OpenStudy (anonymous):

\[\Large\rm \sec ^{4}x \left(\color{lime}{-\sec ^{2}x}\right) +\sec ^{2}x=\sec ^{4}x\] ignore the parentheses, thats subtraction not multiplication but they cancel and sec=sec?

zepdrix (zepdrix):

Good, yes. When you distribute the sec^2x to each term in the brackets, you get THAT (without the brackets). And then a nice cancellation! yay!

zepdrix (zepdrix):

sec^4x=sec^4x

OpenStudy (anonymous):

yep! yay thanks so much :)

zepdrix (zepdrix):

You've probably done enough of these that you're used to the format by now. But just in case, we make sure that we `leave one side of the equation alone, and try to match the other side to it`. So we can't do things like "add this to each side". You probably already knew that :) just making sure though hehe

OpenStudy (anonymous):

yeah haha its a good reminder! thanks again for your help, i really appreciate it.